OpenStudy (photon336):

Find the equation of the circle that passes through (2,3) is tangent to the line 3x-4 = -1

OpenStudy (photon336):

Here's what i've done so far \[(x-h)^{2}+(y-k)^{2} = r \] \[\frac{ d }{ dx }(x-h)^{2}+\frac{ d }{ dx }(y-k)^{2} = 0 \] \[\frac{ dy }{ dx }2(y-k) = -2(x+h)\] \[\frac{ dy }{ dx } = \frac{ -(x+h) }{ (y-k) }\] \[3x-4y = -1 \] \[3x= -1+4y \] \[3x+1 = 4y \] \[\frac{ 3 }{ 4 }x+\frac{ 1 }{ 4 } = y\] since the slope of the tangent line is 3/4 i set this equal to the derivative of the circle \[\frac{ -(x-h) }{ (y-k) } = \frac{ 3 }{ 4 }\]

OpenStudy (photon336):

Now I thought that you just plug in the points x = 2 y = 3 and solve for h and k \[\frac{ -(2-h) }{ (3-k) } = \frac{ 3 }{ 4 } ~h = 5, k = -1 \] \[(x-5)^{2}+(y+1)^{2} = 25 \]

jimthompson5910 (jim_thompson5910):

Your circle goes through (2,3) just fine, but `3x-4y=-1` is not tangent to this circle. http://prntscr.com/b172cg

OpenStudy (photon336):

argh.. I missed something big... the question is actually find the standard equations of the circles that pass through (2,3) and are tangent to both lines (3x-4y) = -1 and 4x+3y = 7..

OpenStudy (photon336):

that's the question word for word. not sure here.

jimthompson5910 (jim_thompson5910):

so the circle would look something like this

jimthompson5910 (jim_thompson5910):

I would set up points P and Q in the form of (x,y) solve each equation for y to get it in terms of x

OpenStudy (photon336):

so I would need two separate equations

OpenStudy (photon336):

So, would one equation be this? \[\frac{ -(x-h) }{ (y-k) } = \frac{ 3 }{ 4 }\] \[-\frac{ 4 }{ 3 }*(x-h)+k = y_{1}\] \[\frac{ -(x-h) }{ (y-k) } = \frac{ -4}{3} = -\frac{ 3 }{ 4 }*(x-h)+k = y_{2}\]

OpenStudy (photon336):

I'm wondering if i'm on the right track with this

jimthompson5910 (jim_thompson5910):

http://assets.openstudy.com/updates/attachments/572d45d3e4b071c0ef3c0e92-jim_thompson5910-1462586209770-circle2.png focus on the purple equation 4x+3y = 7 solve for y to get \[\Large y = \frac{7-4x}{3}\] so if P is some point on this purple equation, and P is allowed to slide anywhere along this line, then point P is \[\Large P = \left(x, \frac{7-4x}{3}\right)\] Agreed?

OpenStudy (photon336):

yes I'm following

jimthompson5910 (jim_thompson5910):

similarly, we solve the red equation 3x - 4y = -1 for y to get \[\Large y = \frac{1+3x}{4}\] making \[\Large Q = \left(x, \frac{1+3x}{4}\right)\] where Q is any point on the red line Agreed?

OpenStudy (photon336):

yep

jimthompson5910 (jim_thompson5910):

so we have some point R that is the center of this circle

OpenStudy (photon336):

following

jimthompson5910 (jim_thompson5910):

we don't know what point R is, but the goal is to make slope of segment RP perpendicular to the purple line slope of segment RQ perpendicular to the red line

jimthompson5910 (jim_thompson5910):

we don't know what R is, but we know its the center of the circle, so R = (h,k)

jimthompson5910 (jim_thompson5910):

`slope of segment RP perpendicular to the purple line` so slope of RP = 3/4 since slope of purple line = -4/3 and you flip the fraction and the sign to get the perpendicular slope

jimthompson5910 (jim_thompson5910):

`slope of segment RQ perpendicular to the red line` so slope of RQ = -4/3 since slope of purple line = 3/4 and you flip the fraction and the sign to get the perpendicular slope

OpenStudy (photon336):

interesting, so you take the slopes of the two lines and take the reciprocal of both.

jimthompson5910 (jim_thompson5910):

are you able to find the slope of RP in general (in terms of x)? hint: slope formula

jimthompson5910 (jim_thompson5910):

sorry I meant in terms of x, h, and k

OpenStudy (photon336):

RP = 3/4 RQ = 4/3

jimthompson5910 (jim_thompson5910):

RQ has negative slope

OpenStudy (photon336):

So this is what I found for the slope taking the derivative of the general formula for the circle. \[-\frac{ (x-h) }{ (y-k) }\]

OpenStudy (photon336):

\[RQ = \frac{ -(x-h) }{ (y-k)} = \frac{ -4 }{ 3 }\] \[RP = \frac{ -(x-h) }{ (y-k)} = \frac{ 3}{ 4 }\]

OpenStudy (photon336):

My next idea is to I guess plug in the points P and Q into the two equations and set them equal to each-other

jimthompson5910 (jim_thompson5910):

\[\Large R = \left(h,k\right) = (x_1,y_1)\] \[\Large P = \left(x, \frac{7-4x}{3}\right) = (x_2,y_2)\] Use the slope formula to find the slope of RP in general (in terms of x,h,k) \[\Large m = \frac{y_2 - y_1}{x_2 - x_1}\] \[\Large m = \frac{\frac{7-4x}{3} - k}{x - h}\] Now plug in the slope of RP (earlier we found it to be 3/4) then isolate the 'x-h' term \[\Large m = \frac{\frac{7-4x}{3} - k}{x - h}\] \[\Large \frac{3}{4} = \frac{\frac{7-4x}{3} - k}{x - h}\] \[\Large 3*(x - h) = 4*\left(\frac{7-4x}{3} - k\right)\] \[\Large x-h = \frac{4*\left(\frac{7-4x}{3} - k\right)}{3}\]

OpenStudy (photon336):

okay, I'm starting to see this a bit better so you would do that for the other point and then with those two equations you can find the value of h, k

jimthompson5910 (jim_thompson5910):

hmm I'm probably making it more complicated than it has to be

OpenStudy (triciaal):